# Super-Resolution works for coastal simulations

### v1

@@ Abstract
Learning fine-scale details of a coastal ocean simulation from a coarse representation is a challenging task. For real-world applications, high-resolution simulations are necessary to advance understanding of many coastal processes, specifically, to predict flooding resulting from tsunamis and storm surges. We propose a Deep Network for Coastal Super-Resolution (DNCSR) for spatiotemporal enhancement to efficiently learn the high-resolution numerical solution. Given images of coastal simulations produced on low-resolution computational meshes using low polynomial order discontinuous Galerkin discretizations and a coarse temporal resolution, the proposed DNCSR learns to produce high-resolution free surface elevation and velocity visualizations in both time and space. To efficiently model the dynamic changes over time and space, we propose grid-aware spatiotemporal attention to project the temporal features to the spatial domain for non-local feature matching. The coordinate information is also utilized via positional encoding. For the final reconstruction, we use the spatiotemporal bilinear operation to interpolate the missing frames and then expand the feature maps to the frequency domain for residual mapping. Besides data-driven losses, the proposed physics-informed loss guarantees gradient consistency and momentum changes. Their combination contributes to the overall 24% improvements in RMSE. To train the proposed model, we propose a large-scale coastal simulation dataset and use it for model optimization and evaluation. Our method shows superior super-resolution quality and fast computation compared to the state-of-the-art methods.
Index Terms:
@@
Coastal simulation, deep neural networks, discontinuous Galerkin method, shallow water equations, super-resolution.
publicationid: pubid: $0$–$00$ 2021 IEEE
I Introduction
@@
The two-dimensional shallow-water equations (SWE) can be utilized to model circulation in the global, regional, and coastal ocean, inland seas, lakes, rivers; they are also frequently employed for atmospheric circulation studies [1]. Currently, SWE-based numerical software packages are the main tool employed in the operational forecast of tsunamis and storm surges [2, 3, 4, 5]. For many coastal ocean applications—flooding simulations are one crucial example—the accuracy of the model results strongly depends on the resolution of the computational mesh and the accuracy of the time discretization. To reliably meet the requirements of accurate prediction of inundation with the purpose of warning and hazard management, mesh resolutions down to and even below 10m in the affected areas [6] are necessary. Even using unstructured scenario-adapted meshes and GPU computing [7, 8], such resolution requirements are computationally challenging—especially in real-time warning systems [9].
@@
Here, $\boldsymbol{q}\coloneqq(U,V)^{T}$ denotes the depth-integrated horizontal velocity, $\xi$ the water elevation above some datum (e.g., the mean sea level), $h_{b}$, the bathymetric depth respective the same datum, and $H\coloneqq h_{b}+\xi$ the total water depth. The other terms have the following meaning: $f_{c}$ the Coriolis coefficient, $g$ the gravitational acceleration, and $\tau_{\textrm{bf}}$ the bottom friction coefficient. Equation (1a) describes the mass conservation and (1b) the conservation of momentum.
The boundary conditions relevant to our scenario are the land boundary condition $\boldsymbol{q}\cdot\boldsymbol{n}=0$ (no water crosses the land boundary; $\boldsymbol{n}$ is a unit normal vector to the boundary) and the open sea boundary condition $\xi=\hat{\xi}$, which prescribes the water elevation at open sea boundaries given as a space- and time-dependent function $\hat{\xi}$.
@@
Since coastal ocean simulations are important and challenging, they have been a focus of many research efforts resulting in numerous numerical and algorithmic advances in recent years. For example, [12] derives an $hp$ adaptive DG discretization, [13, 14] develop hybridized DG methods and suitable time-stepping approaches, whereas [15] proposes an adaptive enriched Galerkin scheme. A multiwavelet approach is explored in [16], while [17] further enhances the concept of adaptivity by omitting computations in the subdomain parts, where no changes take place. Wetting and drying is one of the key processes in coastal regions; it represents a challenge in the numerical solution of the SWE equations since a water height of zero poses significant problems when dividing by $H$ in (1b). Several works [18, 19, 20] propose techniques to address this challenge and the related issues of negative $H$ in DG simulations, which may result in non-physical behavior. Other reasons for potentially unphysical behavior (entropy conditions) are addressed in [21].
@@
The original UTBEST scheme proved very successful and has been further developed and transferred to several code development frameworks such as Matlab/GNU Octave [15], EXASTENCILS code generation framework [22, 23, 24], OpenCL [25, 26], or SYCL [27]. UTBEST also served as a basis for a fully-featured 3D regional ocean model UTBEST3D [28, 29].
This work exploits deep learning as an alternative for high-resolution shallow water simulations. Specifically, we propose a deep learning-based super-resolution network (DNCSR) for coastal simulations. The overall process is shown in Figure 1. We convert the coarse simulations of the target coastal model into three views: the horizontal depth-integrated velocity field $(U,V)$ and the water height ($\xi$). Then, we merge them as RGB image inputs to the proposed DNCSR for super-resolution. The super-resolved outputs can be used in place of fine-resolution simulations e.g. to visualize the results. A major advantage of our approach compared to other ML-based techniques is the fact that the coarse-resolution image originates from the numerical model and conforms to all physical constraints (in particular, the coarse-resolution solution is mass conservative) preventing the reconstructed fine-resolution solution from ’drifting’ away from physical consistency for arbitrarily long simulation runs.
Figure 1: The overall deep learning pipeline for coastal super-resolution. We use the UTBEST model to simulate the coarse solution representation as $U,V,\xi$. We stack them as RGB channels to form a low-resolution (LR) coastal image, which will be taken as input to the proposed DNCSR for spatiotemporal upsampling. The obtained super-resolution (SR) coastal images seamlessly enhance the spatial and temporal details, which will be useful for coastal simulation and visualization.
II Related Works
@@
High-resolution coastal simulation is important for accurately representing complex circulation patterns—particularly relevant in the presence of irregular coastlines and strongly varying topography. However, simulating shallow-water equations at high spatial and temporal resolution requires massive computational efforts. To upsample the low-resolution (LR) scientific data to high-resolution (HR) full-field dynamics is called super-resolution (SR) in image and video processing. Let us revisit related topics in two categories: deep learning for image/video super-resolution and physics-informed super-resolution.
II-A Image/video super-resolution
@@
Image and video super-resolution is an ill-posed problem. One LR image can lead to multiple plausible sharp and clean HR images. The typical solution is to use paired LR-HR data to optimize restoration models via pixel-based loss functions. Since the seminal super-resolution work of Dong et al. [30], many image super-resolution approaches [31, 32, 33, 34, 35, 36, 37] have adopted an end-to-end neural network to learn the regression model for reconstruction. Recently, attention [38] has also been utilized in image super-resolution to involve more pixels for nonlocal pattern exploration. For example, SwinIR [33] proposes to use Swin Transformer [39] for multi-scale nonlocal feature matching, HAN [35] proposes to combine spatial- and channel-wise attention to model the holistic interdependencies among layers and HAT [40] activates more pixels for high-frequency detail reconstruction via window-based and channel attention. To overcome the bottleneck of information loss caused by the deep attention layers, [41] proposes dense-residual connections to mitigate the spatial loss and stabilize the information flow. Generative adversarial networks [42] point to a new direction of photorealistic image super-resolution, where the model learns the manifold of natural images to produce images with pleasant visual quality. SRGAN [32] and ESRGAN [43] are two efficient approaches that can generate super-resolved images with fine details. Following this direction, a lot of works have been exploring denoising diffusion probabilistic model (DDPMs) [44], score-based models [45], and their recent variations [46, 47, 48, 49, 50] for generative image super-resolution. SR3 [31] and StableSR [34] are two representative approaches that achieve photorealistic image reconstruction. However, they also suffer from high computational costs and slow inference.
@@
II-B Physics-informed super-resolution
@@
Thanks to the great success in image and video super-resolution, considerable studies exist about the interplay between physics and machine learning. The main trend is to approximate the HR data representation based on potentially noisy and under-resolved simulations, like smoke [55], climate [56, 57], and chemistry [58, 59]. Most deep learning-based SR for scientific data can be categorized into two groups: spatial super-resolution and temporal super-resolution. For the former, it is similar to image super-resolution where, given the LR data representation, the network should produce fine-grained data representation. For example, Fukami et al. [60] use the SRCNN [30] model to upsample 2D laminar cylinder flow. MeshfreeFlowNet [61] is proposed to reconstruct the turbulent flow in the Rayleigh-Benard problem via a UNet structure. Given the challenge in the laminar finite-rate-chemistry flows, [62] proposes to use PIESRGAN to estimate the high-resolution flow. Using subgrid turbulent flow models, the idea is to extend ESRGAN [43] to the 3D space. Two physics-informed losses, gradient loss and continuity loss, contribute to the gradient and total mass changes. PINNSR proposes to use RRDB blocks [63] to build a GAN-like network to simulate the fine-grid Rayleigh-Taylor instability. The physical loss is used to govern the advection-diffusion process. PhySR [64] proposes to use a physics-informed network to learn temporal super-resolution. The key idea is to use ConvLSTM [65] to learn temporal refinement and dynamic evolution on LR features. Stacked residual blocks are used to learn pixel reconstruction. PhySRNet [66] proposes an unsupervised learning approach to approximate the high-resolution counterparts without requiring labeled data.
@@
Similarly, Gao et al. [67] also propose using conservation laws and boundary conditions of fluid flows so that the model can be optimized in a self-supervision manner. Teufel et al. [68] propose to predict fine-grid regional climate simulations via an FSR network, which can learn temporal interpolation via flow warping. [69] modifies the diffusion model to simulate the computational fluid dynamics data. It is trained only on high-resolution to learn multi-scale downsampling scenarios, resulting in a robust super-resolution model for 2D turbulent flow estimation. Distinct from existing methods, ours focuses on spatiotemporal super-resolution and explicitly explores the multidimensional signal via a spatiotemporal attention module. We also explore physics-informed loss to govern the optimization process and show significant improvements compared to others.
@@
It is worth noticing that applying deep neural networks to the real physical worldmust produce results which guarantee conservation of key physical quantities such as mass or energy over long simulation time intervals. For instance, given paired LR and HR image pairs, the sum of values at local HR pixels needs to match with corresponding LR pixels. In deep neural networks, the common solution is using physics-informed losses. For example, Beucler et al. [70] propose soft penalties on the loss terms to emulate the cloud process, so that the network follows the conservation of enthalpy, the conservation of mass, terrestrial radiation, and solar radiation. The high intrinsic uncertainty of the climate system makes it difficult for neural networks to predict long-term simulation, Harder et al. [71] propose a correction layer attached to the neural network and train it via physical data generated by the aerosol microphysics model. It guarantees perfect mass conservation and significant speedup. To better preserve the physical quantities, [72] proposes multiple local average pooling constraints to supervise the training optimization, hence it can zero violation of the physics, produce no negative pixels, and preserve mass up to numerical precision. In contrast to all the above methods, we propose to use a numerical base method that guarantees the conservation of mass and momentum and conducts super-resolution concerning the results of this method. Thus, we guarantee that mass and momentum are conserved over time and enable reliable long-term simulations.
III Approach
@@
This section introduces our workflow, which includes the coastal ocean simulation and the proposed deep neural network for coastal super-resolution.
III-A Problem setup for coastal simulation
@@
Given a specific coastal region that we want to model, we first use our DG-based SWE model code for hydrodynamics simulation. Each coastal region is simulated with varying polynomial approximation orders, mesh resolutions, and time steps to capture different levels of detail and accuracy. The simulations are then rendered as images to facilitate the application of computer vision techniques in neural network training, leveraging the spatial structure of the data. Specifically, we convert the simulations into three views: the horizontal depth-integrated velocity field ($U$,$V$) and the water elevation ($\xi$), stacking them as RGB channels to create a comprehensive visual representation of the fluid dynamics. For each resolution, we employ three different modes of grid interpolation to examine the impact of interpolation methods on the simulation accuracy and neural network training effectiveness. These simulations at varying spatial and temporal resolutions are utilized for neural network training to enhance the model’s ability to generalize across different scales and time frames.
III-B Proposed DNCSR for coastal super-resolution

### v2

# Downscaling Neural Network for Coastal Simulations

@@ Abstract
Learning the fine-scale details of a coastal ocean simulation from a coarse representation is a challenging task. For real-world applications, high-resolution simulations are necessary to advance understanding of many coastal processes, specifically, to predict flooding resulting from tsunamis and storm surges. We propose a Downscaling Neural Network for Coastal Simulation (DNNCS) for spatiotemporal enhancement to efficiently learn the high-resolution numerical solution. Given images of coastal simulations produced on low-resolution computational meshes using low polynomial order discontinuous Galerkin discretizations and a coarse temporal resolution, the proposed DNNCS learns to produce high-resolution free surface elevation and velocity visualizations in both time and space. To efficiently model the dynamic changes over time and space, we propose grid-aware spatiotemporal attention to project the temporal features to the spatial domain for non-local feature matching. The coordinate information is also utilized via positional encoding. For the final reconstruction, we use the spatiotemporal bilinear operation to interpolate the missing frames and then expand the feature maps to the frequency domain for residual mapping. Besides data-driven losses, the proposed physics-informed loss guarantees gradient consistency and momentum changes. Their combination contributes to the overall 24% improvements in Root Mean Square Error (RMSE). To train the proposed model, we propose a novel coastal simulation dataset and use it for model optimization and evaluation. Our method shows superior downscaling quality and fast computation compared to the state-of-the-art methods.
keywords:
@@
Coastal simulation , deep neural networks , discontinuous Galerkin method , shallow water equations , downscaling.
journal: Advances in Water Resources
[…]
[a]organization=Department of Computational Engineering, Lappeenranta-Lahti University of Technology LUT, city=Lahti, country=Finland
[…]
[b]organization=University of Bayreuth, city=Bayreuth, country=Germany
[…]
[c]organization=Saarland University, city=Saarbrücken, country=Germany
1 Introduction
@@
The two-dimensional shallow-water equations (SWE) can be utilized to model circulation in the global, regional, and coastal ocean, inland seas, lakes, and rivers; they are also frequently employed for atmospheric circulation studies [72]. Currently, SWE-based numerical software packages are the main tool employed in the operational forecast of tsunamis and storm surges [16, 5, 79, 17]. For many coastal ocean applications—flooding simulations are one crucial example—the accuracy of the model results strongly depends on the resolution of the computational mesh and the accuracy of the time discretization. To reliably meet the requirements of accurate prediction of inundation with the purpose of warning and hazard management, mesh resolutions down to and even below 10m in the affected areas [10] are necessary. Even using unstructured scenario-adapted meshes and GPU computing [55, 63], such resolution requirements are computationally challenging—especially in real-time warning systems [56].
Since coastal ocean simulations are important and challenging, they have been a focus of many research efforts resulting in numerous numerical and algorithmic advances in recent years. While many established software packages rely on Finite Volume of Finite Difference methods, the discontinuous Galerkin (DG) and other related discretization techniques (such as enriched Galerkin or hybridized DG methods) emerged in the last two decades as strong candidates, especially for unstructured mesh and adaptive simulations. For example, [24] derives an $hp$ adaptive DG discretization, [15, 62] develop hybridized DG methods and suitable time-stepping approaches, whereas [39] proposes an adaptive enriched Galerkin scheme. A multiwavelet approach is explored in [32], while [36] further enhances the concept of adaptivity by omitting computations in the subdomain parts, where no changes take place.
Recently, there have been rapid developments of deep learning based super-resolution in image and video processing [85, 75, 76, 52]. The ill-posed problems in image processing are considered to be similar to the downscaling problem in scientific modeling, like fluid dynamic simulation [66], climate downscaling [7], and smoke simulation [8]. However, there are no or very few works on applying deep learning approaches for coastal ocean downscaling. This work exploits deep learning as an alternative for high-resolution (HR) shallow water models. Specifically, we propose a Downscaling Neural Network for Coastal Simulations (DNNCS) to downscale the results of low-resolution (LR) coastal simulations to obtain HR results. The overall process is shown in Figure 1. We convert the coarse simulations of the target coastal model into three views: the horizontal depth-integrated horizontal velocity field $(U,V)$ and the water height ($\xi$). Then, we merge them as RGB image inputs to the proposed DNNCS for downscaling. The super-resolved outputs can be used in place of fine-resolution simulations, e.g., to visualize the results. Besides the fast computation for HR shallow water simulations, another major advantage of our approach compared to other ML-based techniques is the fact that the coarse-resolution image originates from the numerical model and conforms to all physical constraints (in particular, the coarse-resolution solution is mass conservative), preventing the reconstructed fine-resolution solution from ’drifting’ away from physical consistency for arbitrarily long simulation runs.
Figure 1: The overall deep learning pipeline for coastal ocean downscaling. We use the numerical model to simulate the coarse solution representation as $U,V,\xi$. We stack them as RGB channels to form a low-resolution (LR) image, which will be taken as input to the proposed DNNCS for spatiotemporal upsampling. The obtained downscaling coastal images seamlessly enhance the spatial and temporal details, which will be useful for coastal ocean simulation and visualization.
To summarize, our main contributions are as follows:
  • 1.
    We present a novel DNNCS for coastal ocean simulation, capable of efficiently generating high-resolution data and reducing the computational burden associated with PDE-based physical modeling.
  • 2.
    We introduce a spatio-temporal attention mechanism to jointly learn spatial and temporal downscaling, incorporating coordinate information as positional signals for grid-aware reconstruction.
  • 3.
    To ensure physical consistency, we incorporate physics-informed loss functions as constraints, guiding the model to produce reliable high-resolution outputs.
  • 4.
    We construct a new coastal simulation dataset featuring multiple levels of resolution to support effective model training and validation.
2 Governing equations and simulation software
@@
Here, $\boldsymbol{q}\coloneqq(U,V)^{T}$ denotes the depth-integrated horizontal velocity, $\xi$ the water elevation above some datum (e.g., the mean sea level), $h_{b}$ the bathymetric depth respective the same datum, and $H\coloneqq h_{b}+\xi$ the total water depth. The other terms have the following meaning: $f_{c}$ is the Coriolis coefficient, $g$ is the gravitational acceleration, and $\tau_{\textrm{bf}}$ is the bottom friction coefficient.
The boundary conditions relevant to our scenario are the land boundary condition $\boldsymbol{q}\cdot\boldsymbol{n}=0$ and the open sea boundary condition $\xi=\hat{\xi}$, which prescribes the tidal elevation at open sea boundaries given as a space- and time-dependent function $\hat{\xi}$.
@@
The original UTBEST scheme proved very successful and has been further developed and transferred to several code development frameworks such as Matlab/GNU Octave [39], EXASTENCILS code generation framework [26, 4, 27, 25], OpenCL [44, 28], or SYCL [18, 19]. UTBEST also served as a basis for a fully-featured 3D regional ocean model UTBEST3D [22, 2].
3 Related Works
@@
High-resolution coastal ocean simulation is important for accurately representing complex circulation—particularly relevant in the presence of irregular coastlines and strongly varying topography. However, simulating shallow-water equations at high spatial and temporal resolution requires massive computational efforts. To upsample the low-resolution (LR) scientific data to high-resolution (HR) full-field dynamics is called super-resolution (SR) in image and video processing, and downscaling in climate and weather prediction. Let us revisit related topics in two categories: deep learning for image/video super-resolution and physics-informed downscaling.
3.1 Learning based super-resolution and downscaling
@@
Image and video super-resolution is an ill-posed problem. One LR image can lead to multiple plausible sharp and clean HR images. The typical solution is to use paired LR-HR data to optimize restoration models via pixel-based loss functions. Since the seminal super-resolution work of Dong et al. [23], many image super-resolution approaches [54, 53, 47, 50, 75, 58, 71, 43] have adopted an end-to-end neural network to learn the regression model for reconstruction. Recently, attention [73] has also been utilized in image super-resolution to involve more pixels for nonlocal pattern exploration. For example, SwinIR [50] proposes to use Swin Transformer [51] for multi-scale nonlocal feature matching, HAN [58] proposes to combine spatial- and channel-wise attention to model the holistic interdependencies among layers and HAT [20] activates more pixels for high-frequency detail reconstruction via window-based and channel attention. To overcome the bottleneck of information loss caused by the deep attention layers, [41] proposes dense-residual connections to mitigate the spatial loss and stabilize the information flow. Generative adversarial networks [34] point to a new direction of photorealistic image super-resolution, where the model learns the manifold of natural images to produce images with pleasant visual quality. SRGAN [47] and ESRGAN [77] are two efficient approaches that can generate super-resolved images with fine details. Following this direction, a lot of works have been exploring denoising diffusion probabilistic model (DDPMs) [67], score-based models [68], and their recent variations [69, 9, 65, 40, 60] for generative image super-resolution. SR3 [61] and StableSR [75] are two representative approaches that achieve photorealistic image reconstruction. However, they also suffer from high computational costs and slow inference.
@@
With these advances in deep learning-based super-resolution (SR), its application to climate and weather prediction has gained significant attention, namely climate/weather downscaling. That is, downscaling the grid size for ocean and weather simulation so we can get fine-resolution visualization. For example, it can be used in the sea surface temperature (SST) for high-resolution air-sea interaction study [29, 45]. It can also find the links between coarse sea surface height (SSH) and high-resolution SST fields [70, 82], which can help to understand ocean circulation and sea surface topography. Several works [3, 12, 38, 57] have explored downscaling spatial resolution from 50$\sim$100 km to 1$\sim$25 km for climate data, and improving temporal resolution to hourly data. This enables more accurate regional infrastructure planning, resource adequacy evaluation, and risk assessment.
3.2 Physics-informed downscaling/super-resolution
@@
Thanks to the great success in image and video super-resolution, considerable studies exist about the interplay between physics and machine learning. The main trend is to approximate the HR data representation based on potentially noisy and under-resolved simulations, like smoke [8], climate [3, 13], and chemistry [33, 74]. Most deep learning-based SR for scientific data can be categorized into two groups: spatial downscaling and temporal downscaling. For the former, it is similar to image super-resolution where, given the LR data representation, the network should produce fine-grained data representation. For example, Fukami et al. [30] use the SRCNN [23] model to upsample 2D laminar cylinder flow. MeshfreeFlowNet [42] is proposed to reconstruct the turbulent flow in the Rayleigh-Benard problem via a UNet structure. Given the challenge in the laminar finite-rate-chemistry flows, [14] proposes to use PIESRGAN to estimate the high-resolution flow. Using subgrid turbulent flow models, the idea is to extend ESRGAN [77] to the 3D space. Two physics-informed losses, gradient loss and continuity loss, contribute to the gradient and total mass changes. PINNSR proposes to use RRDB blocks [85] to build a GAN-like network to simulate the fine-grid Rayleigh-Taylor instability. The physical loss is used to govern the advection-diffusion process. PhySR [59] proposes to use a physics-informed network to learn temporal downscaling. The key idea is to use ConvLSTM [64] to learn temporal refinement and dynamic evolution on LR features. Stacked residual blocks are used to learn pixel reconstruction. PhySRNet [6] proposes an unsupervised learning approach to approximate the high-resolution counterparts without requiring labeled data.
@@
Similarly, Gao et al. [31] also propose using conservation laws and boundary conditions of fluid flows so that the model can be optimized in a self-supervised manner. Teufel et al. [7] propose to predict fine-grid regional climate simulations via an FSR network, which can learn temporal interpolation via flow warping. [66] modifies the diffusion model to simulate the computational fluid dynamics data. It is trained only on high-resolution to learn multi-scale downsampling scenarios, resulting in a robust downscaling model for 2D turbulent flow estimation. Distinct from existing methods, ours focuses on spatiotemporal downscaling and explicitly explores the multidimensional signal via a spatiotemporal attention module. We also explore physics-informed loss to govern the optimization process and show significant improvements compared to others.
@@
It is worth noticing that applying deep neural networks to the real physical world must produce results that guarantee conservation of key physical quantities such as mass or energy over long simulation time intervals. For instance, given paired LR and HR image pairs, the sum of values at local HR pixels needs to match with corresponding LR pixels. In deep neural networks, the common solution is using physics-informed losses. For example, Beucler et al. [11] propose soft penalties on the loss terms to emulate the cloud process, so that the network follows the conservation of enthalpy, the conservation of mass, terrestrial radiation, and solar radiation. The high intrinsic uncertainty of the climate system makes it difficult for neural networks to predict long-term simulations. Harder et al. [37] propose a correction layer attached to the neural network and train it via physical data generated by the aerosol microphysics model. It guarantees perfect mass conservation and significant speedup. To better preserve the physical quantities, [38] proposes multiple local average pooling constraints to supervise the training optimization, hence it can have zero violation of the physics, produce no negative pixels, and preserve mass up to numerical precision. In contrast to all the above methods, we propose to use a numerical base method that guarantees the conservation of mass and momentum and conducts downscaling concerning the results of this method. Thus, we guarantee that mass and momentum are conserved over time and enable reliable long-term simulations.
4 Approach
@@
This section introduces our workflow, which includes the coastal ocean simulation and the proposed downscaling neural network.
4.1 Problem setup for coastal simulation
@@
Given a specific coastal region that we want to model, we first use our DG-based SWE model code for hydrodynamics simulation. Each coastal region is simulated with varying polynomial approximation orders, mesh resolutions, and time steps to capture different levels of detail and accuracy. The simulations are then rendered as images to facilitate the neural network training, leveraging the spatial structure of the data. Specifically, we convert the simulations into three views: the horizontal depth-integrated velocity field ($U$,$V$) and the water elevation ($\xi$), stacking them as RGB channels to create a comprehensive visual representation of the ocean dynamics. For each resolution, we employ three different modes of grid interpolation to examine the impact of interpolation methods on the simulation accuracy and neural network training effectiveness. These simulations at varying spatial and temporal resolutions are utilized for neural network training to enhance the model’s ability to generalize across different scales and time frames.
4.2 Proposed DNNCS for coastal downscaling